What is the volume of the region under the surface z=xy that lies over the triangle with vertices at the origin, (3, 0), and (0, 6)?
1
Expert's answer
2020-10-25T12:22:17-0400

The equation of line joining the points "(3,0)" and "(0,6)" is given by,


"y-0=\frac{6-0}{0-3}(x-3)"


"y=-2(x-3)"


"y=6-2x"


Thee triangular region with vertices at origin, "(3,0)" and "(0,6)" is as shown in the figure below:





Now, the volume of the region under the surface "z=xy" over the triangular region is evaluated as,


"V=\iint_{R}zdA"


"=\int_{0}^{3}\int_{0}^{6-2x}xydydx"


"=\int_{0}^{3}x[\frac{y^2}{2}]_{0}^{6-2x}dx"


"=\frac{1}{2}\int_{0}^{3}x(6-2x)^2dx"


"=\frac{1}{2}\int_{0}^{3}x(36+4x^2-24x)dx"


"=\int_{0}^{3}x(18+2x^2-12x)dx"


"=\int_{0}^{3}(18x+2x^3-12x^2)dx"


"=[18(\frac{x^2}{2})+2(\frac{x^4}{4})-12(\frac{x^3}{3})]_{0}^{3}"


"=9(9-0)+\frac{1}{2}(81-0)-4(27-0)"


"=81+\frac{81}{2}-108"


"=\frac{81}{2}-27"


"=\frac{27}{2}"


Therefore, the volume of region under the surface "z=xy" is "V=\frac{27}{2}" cubic units.